OpenStudy (bbb911):

____4x^4____ / _____2x^3_____ x^2 - 1 x^2 - 2x + 1 Divide & Simplify

OpenStudy (anonymous):

First we factor x^2 - 1 = x^2 - 1^2 = (x -1) * (x + 1) okay ?

OpenStudy (bbb911):

okay

OpenStudy (anonymous):

& also (x^2 - 2x + 1) in perfect binomial as (x - 1)^2 right ?

OpenStudy (bbb911):

ok

OpenStudy (anonymous):



OpenStudy (bbb911):

okay so you factored the bottom first

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):